This Question: 1 pt 4 of 5 (0 comple Find the real solutions of the equation. 14x + 51-5 Select the correct choice below and il in any answer boxes within your choice O A. The solution set is O B. There is no real solution. (Simplify your answer. Use a comma to separate answers as needed) Click to select and enter your answer(s) ns n 24/7
Answer #1

Solution:

\small \left | x \right |=\left\{\begin{matrix} x &\: \: \: x\geq 0\\ -x &\: \: \: x< 0 \end{matrix}\right.

\small \left | 4x+5 \right |=5

\small \therefore \pm \left ( 4x+5 \right )=5

\small \therefore 4x+5=5 \: \: \: \: \: \: \:,\: \: \: \: \: 4x+5=-5

\small \therefore4x=0 \: \: \: \: \: \: \:,\: \: \: \: \: 4x=-5-5

\small \therefore\: x=0 \: \: \: \: \: \: \:,\: \: \: \: \: 4x=-10

\small \therefore\: x=0 \: \: \: \: \: \: \:,\: \: \: \: \: x=\frac{-10}{4}=\frac{-5}{2}

The solution set is \small {\color{Magenta} \left \{ 0\: ,\: \frac{-5}{2} \right \}} .